A ship moves due North at 100km/h on a river flowing due East at 25km/h. Calculate the magnitude of the resultant velocity of the ship and it direction.


1
Expert's answer
2021-09-08T16:52:46-0400

We can find the magnitude of the resultant velocity of the ship from the Pythagorean theorem:


"v_{res}=\sqrt{v_{river}^2+v_{ship}^2}=\sqrt{(25\ \dfrac{km}{h})^2+(100\ \dfrac{km}{h})^2}=103\ \dfrac{km}{h}."

We can find the direction of the resultant velocity from the geometry:


"\theta=tan^{-1}(\dfrac{v_{river}}{v_{ship}}),""\theta=tan^{-1}(\dfrac{25\ \dfrac{km}{h}}{100\ \dfrac{km}{h}})=14^{\circ}\ N\ of\ E."